Overhead ABC

AJ Accounting performs two types of services, Audit and Tax. AJ's overhead costs consist of computer support, $200,000; and legal support, $100,000. Information on thetwo services is:
Audit Tax
Direct Labor $50,000 $100,000
CPU minutes 40,000 10,000
Legal hours used 200 800
Overhead applied to audit services using activity-based costing is?
a. $100,000
b. $120,000
c. $180,000
d. $200,000
Answer #1
Overhead applied to audit services using activity-based costing is?

c. $180,000

Computer support 200,000 is divided in the ratio of 40,000 and 10,000 i.e 160,000
legal support 100,000 is divided in the ratio of 200 and 800 i.e 20,000
total is 180,000
answered by: mla
Answer #2
Total CPU minuts=40,000+10,000=50,000
Overhead applied to audit -CPU min($200,000/50,000)*40,000=$4/min*40,000=$160,000
Overhead applied to aidut-Legal hoours used($100,000/(200+800))*200=$20,000
Total overhead applied to audit using activity costing=$160,000+$20,000=$180,000
c.
answered by: torrey